OpenStudy (anonymous):

What is the probability of hitting each target if it is assumed that a player’s throw will always hit the board? Hint: The probability of hitting each target is the area of the target hole divided by the total area of the target. Take the value of π to be 3.14. @zepdrix

OpenStudy (mathmate):

Example:  Area of board = \(\pi r^2=\pi (3^2)=9\pi\) Area of target = \(\pi r^2=\pi (0.5^2)=0.25\pi\) Assuming the hits are random, Probability of hitting target = \(0.25\pi / 9\pi = 1/36\)